OpenStudy (anonymous):

determine whether the equation 4x^2 - 9y^2 = 36 defines as a function of x

OpenStudy (anonymous):

In that question... just replace (-x) or (-y) to the equation.. The one that you can get the same answer will be its function... -- am i right @hartnn ?

OpenStudy (sweetburger):

the equation is a function of x

OpenStudy (anonymous):

4x² + 9y² = 36 Subtract 4x² from both sides: 9y² = 36 - 4x² Divide both side by 9: y² = {36 - 4x²}/9 Factor out (4/9) on the right: y² = {4/9}{9 - x²} Take the square root of both sides: y = (2/3)sqrt{9 - x²} y = -(2/3)sqrt{9 - x²} Check by starting with the original equation and substituting both values for y: 4x² + 9y² = 4x² + 9{(2/3)sqrt{9 - x²}}² = 4x² + 9{(4/9){9 - x²} = 4x² + 4{9} - 4x² = 36 4x² + 9y² = 4x² + 9{-(2/3)sqrt{9 - x²}}² = 4x² + 9{(4/9){9 - x²} = 4x² + 4{9} - 4x² = 36

OpenStudy (anonymous):

9y² = 36 - 4x² = 4(9 - x²) y² = 4/9*(9 - x²) √y² =√ [4/9*(9 - x²)] y = ±2/3√(9 - x²) I don't know where you got that ±√(x - 1) Your equation is an ellipse. Divide all by 36 x²/9 + y²/4 = 1 the center is (0, 0) the major axis is "x" with a length of 3 while the minor axis is y with a length of 2.